### v1

# RefDrone: A Challenging Benchmark for Referring Expression Comprehension in Drone Scenes

Abstract
Drones have become prevalent robotic platforms with diverse applications, showing significant potential in Embodied Artificial Intelligence (Embodied AI). Referring Expression Comprehension (REC) enables drones to locate objects based on natural language expressions, a crucial capability for Embodied AI. Despite advances in REC for ground-level scenes, aerial views introduce unique challenges including varying viewpoints, occlusions and scale variations. To address this gap, we introduce RefDrone, a REC benchmark for drone scenes. RefDrone reveals three key challenges in REC: 1) multi-scale and small-scale target detection; 2) multi-target and no-target samples; 3) complex environment with rich contextual expressions. To efficiently construct this dataset, we develop RDAgent (referring drone annotation framework with multi-agent system), a semi-automated annotation tool for REC tasks. RDAgent ensures high-quality contextual expressions and reduces annotation cost. Furthermore, we propose Number GroundingDINO (NGDINO), a novel method designed to handle multi-target and no-target cases. NGDINO explicitly learns and utilizes the number of objects referred to in the expression. Comprehensive experiments with state-of-the-art REC methods demonstrate that NGDINO achieves superior performance on both the proposed RefDrone and the existing gRefCOCO datasets. The dataset and code will be publicly at https://github.com/sunzc-sunny/refdrone.
1 Introduction
Figure 1: Examples of the various challenges in RefDrone dataset.
Table 1: Comparison of referring expression comprehension datasets: RefCOCO/+/g [45, 29], gRefCOCO [22], D3 [42], RIS-CQ [13], RSVG [47], and RefDrone (ours). Avg. object: average number of objects per expression. Avg. length: words per expression. No target: expressions without referred objects. Expression type: expression generation method. Small target: small-scale target detection challenge.
RefCOCO/+/g [45, 29] gRefCOCO [22] D3 [42] RIS-CQ [13] RSVG [47] RefDrone Image source COCO [20] COCO [20] COCO [20] VG [16]+COCO [20] DIOR [18] VisDrone [49] Avg. object 1.0 1.4 1.3 3.6 2.2 3.8 Avg. length 3.6/3.5/8.4 4.9 6.3 13.2 7.5 9.0 No target Expression type Manual Manual Manual LLM (GPT3.5-t) Templated LMM (GPT4-o) Small target
Drones/UAVs have become increasingly popular in our daily lives, serving both personal and professional purposes, such as entertainment, package delivery, traffic surveillance, and emergency rescue. Their ability to move freely and their broad observation capabilities make them important platforms for Embodied AI [6, 27, 17, 8, 25]. A crucial capability in Embodied AI is Referring Expression Comprehension (REC) [35, 40, 2, 39], which serves as a critical bridge between natural language understanding and visual perception. REC requires drones to localize specific objects in images based on natural language expressions. However, existing REC datasets primarily focus on ground-level perspectives, such as the RefCOCO [45, 29] dataset. The application of REC in drone-based scenarios presents unique challenges, including extreme viewpoint variations, occlusions, and scale variations across objects.
In this work, we introduce RefDrone, a challenging benchmark designed for REC in drone scenes. The RefDrone dataset contains 17,900 referring expressions across 8,536 images, comprising 63,679 object instances. As illustrated in Figure 1, RefDrone poses three key challenges that distinguish it from existing datasets: (1) multi-scale and small-scale target detection, with 31% small objects and 14% large objects; (2) multi-target and no-target samples, where expressions can refer to any number of objects from 0 to 242; and (3) complex environment with rich contextual expressions. The environmental complexity manifests in varied viewpoints, diverse lighting conditions, and complex backgrounds. The contextual richness is reflected in expressions that describe spatial relations, attributes, and interactive relationships.
To efficiently construct this comprehensive dataset, we develop RDAgent (referring drone annotation framework with multi-agent system), a semi-automated annotation pipeline. RDAgent restructures traditional annotation workflows into an interactive system of multiple agents and human annotators. In this framework, agents with diverse roles collaborate through feedback loops to generate and verify annotations, while human involvement is minimized to quality control and minor adjustments. RDAgent significantly reduces annotation costs while maintaining high-quality standards for complex expressions, and it can be extended to generate annotations for other REC tasks.
Furthermore, we propose a novel method called Number GroundingDINO (NGDINO). Our key insight is that utilizing the number information of referred objects enhances the handling of multi-target and no-target samples. NGDINO comprises three key components: (1) a number prediction head that estimates target object quantities, (2) a set of learnable number-queries encoding numerical patterns across different quantity samples, and (3) a number cross-attention module that integrates number queries with detection queries. Extensive experiments on both our drone-specific RefDrone benchmark and the general-domain gRefCOCO [22] dataset reveal that leveraging the number information greatly improves performance on multi-target and no-target samples.
In summary, our contributions are listed as follows:
  1. 1.
    RefDrone Benchmark: The first comprehensive benchmark for referring expression comprehension in drone scenes. RefDrone poses three key challenges and provides comprehensive baseline evaluations across both specialized models and large multimodal models.
  2. 2.
    RDAgent Annotation Framework: A novel semi-automated annotation framework employs a multi-agent system. RDAgent reduces annotation costs while ensuring high-quality referring expressions.
  3. 3.
    NGDINO Method: A novel method specifically designed to handle multi-target and no-target samples through three components: a number prediction head, learnable number-queries, and number cross-attention. NGDINO achieves superior performance on both RefDrone and gRefCOCO datasets.
2 Related works
2.1 Referring expression understanding datasets
Referring expression understanding identifies specific regions within images or videos using natural language expressions. Two primary subtasks are Referring Expression Comprehension (REC) and Referring Expression Segmentation (RES). REC outputs detection bounding boxes, while RES outputs segmentation masks. Early datasets such as ReferIt [15] and RefCOCO [45] are pioneering but are limited to single-target expressions and simple expressions. Subsequently, datasets [22, 37, 13, 42] introduce more challenges. gRefCOCO [22] supports multi-target expressions. OV-VG [37] enables an open-vocabulary setting. D3 [42] and RIS-CQ [13] introduce more complex expressions. Additionally, domain-specific datasets [47, 30, 32] emerge to address unique applications. RSVG [47] focuses on remote sensing, RAVAR [30] targets video action recognition, and RIO [32] addresses affordance detection. However, drone scenes remain unexplored in REC tasks, which motivates our introduction of the RefDrone benchmark.
Recent advances in dataset creation have been significantly influenced by large language models (LLMs). Notable works such as LLaVA [24] and Ferret [44] show the effectiveness of LLMs in generating instruction tuning data. Similarly, RIS-CQ [13] and RIO [32] employ LLMs to generate complex referring expressions. However, these approaches have two limitations. First, LLMs are used mainly as text generators, potentially producing expressions without proper grounding in visual context. Second, the lack of iterative feedback mechanisms in the annotation pipelines may degrade annotation quality.
Figure 2: The overview of the RefDrone annotation process with RDAgent. Multiple specialized agents collaborate both with each other and human annotators through iterative feedback loops to generate high-quality annotations.
2.2 Referring expression comprehension methods
Referring expression comprehension methods can be broadly categorized into specialist models and large multimodal models (LMMs). Specialist models include two-stage and one-stage methods. Two-stage methods [12, 23, 10, 36, 9] typically approach REC as a ranking task, first generating region proposals and then ranking them based on language input. However, they are often criticized for slow inference speeds. In contrast, one-stage methods [14, 19, 26, 7, 43, 28] directly predict target regions guided by language. These approaches leverage transformer architectures to enable cross-modal interactions between visual and textual features. Among these methods, GroundingDINO (GDINO) [28] has gained widespread attention for its impressive results in both open-set object detection and REC tasks. However, GDINO lacks explicit mechanisms to handle multi-target or no-target scenarios.
Recent studies have increasingly evaluated LMMs [24, 44, 3, 5, 1, 31, 38, 21, 48, 41] on REC tasks to assess their visual-language understanding capabilities. These models leverage large-scale referring instruction tuning data to achieve remarkable performance. Despite their broad capabilities, LMMs consistently struggle with small object detection, primarily due to their inherent input resolution constraints. This constraint forces image downsampling, resulting in the loss of fine-grained visual details.
3 RefDrone benchmark
3.1 Data source
The RefDrone dataset builds upon VisDrone2019-DET [49], a drone-captured dataset for object detection. The source images are collected across multiple scenarios, illumination conditions, and flying altitudes. To ensure meaningful visual content, we implement two filtering criteria. At the image level, we exclude images containing less than 3 objects. At the object level, we exclude objects with bounding box areas below 64 pixels. The original VisDrone2019-DET annotations provide object categories and bounding box coordinates. We convert the bounding box to a normalized center point (range 0-1). This approach reduces the input context for LMMs while preserving spatial information.
3.2 RDAgent for semi-automated annotation
RDAgent (referring drone annotation framework with multi-agent system) is a semi-automated annotation framework that enhances traditional annotation pipelines by integrating LMMs with human annotators. Central to RDAgent are multiple LMM agents, specifically GPT-4o, each perform specialized tasks via distinct system prompts. As shown in Figure 2, RDAgent employs five structured steps.
Step 1: scene understanding. A GPT-4o-based captioning agent generates three diverse textual descriptions per image. These captions provide contextual perspectives (e.g., object relationships, spatial layouts) to anchor subsequent referring expression generation.
Figure 3: Object size distribution analysis. (a) Object size distribution in RefDrone dataset (small: $<32^{2}=1024$ pixels, normal: 1024 to 9216 pixels, large: $>96^{2}=9216$ pixels). (b) Object size histograms in RefDrone and gRefCOCO [22] datasets.
Step 2: color categorization. Color attributes serve as crucial discriminative features in referring expressions. Color attributes are extracted using a hybrid pipeline: (1) a WideResNet-101 [46] model trained on HSV color space-based labels, refined via manual validation, and (2) a dedicated LMM agent that verifies predictions to mitigate illumination- or occlusion-induced errors.
Step 3: expression generation. To address the complexity in referring expression generation, we reformulate the referring expression generation task as an object grouping problem. The target of the agent is to group semantically related objects and to provide appropriate reasons for each group. The reasons serve as the referring expressions. Furthermore, a dynamic feedback loop triggers color categorization (Step 2) when novel colors are detected, ensuring color attribute consistency.
Step 4: quality evaluation. An evaluator agent assesses the annotation semantic accuracy and referential uniqueness of object-expression pairs, categorizing annotations as:
  • “Yes.” The annotation is accurate and unique to the described objects.
  • “No.” The annotation is inaccurate or non-unique, with a detailed explanation of the discrepancy.
Annotations marked “Yes” proceed to Step 5. Annotations marked “No” activate specific feedback: expressions with semantic issues are returned to Step 3 (expression generation), while color-related inaccuracies are sent to Step 2 (color categorization) for refinement.
Step 5: human verification. Human annotators review annotation outputs in three tiers:
  • Direct acceptance. Annotations satisfying all quality criteria are approved for the final dataset.
  • Refinement required. Annotations with minor errors are corrected through human editing.
  • Significant issues. Annotations with major errors or unclear descriptions are marked for complete reworking.
Annotations with significant issues enter a feedback loop, returning to Step 4 (quality evaluation). If an annotation repeatedly fails to meet standards, it becomes a no-target sample. These feedback refinement loops ensure that all annotations meet our quality standards and that no-target expressions still relate to the image content.
Each step utilizes LMMs through in-context learning with manually designed prompts and examples (see Supplementary Material). RDAgent requires only \$0.0539 pre expression in GPT4-o API cost and reduces human annotation effort by 88% compared to fully manual annotations. This cost-performance balance makes the framework scalable for large-scale datasets.
Figure 4: Object number distribution per expression in gRefCOCO [22] and RefDrone datasets.
Figure 5: Word frequency visualization in RefDrone dataset.
3.3 Dataset analysis
The RefDrone dataset contains 17,900 referring expressions across 8,536 images, comprising 63,679 object instances in 10 categories. We maintain the original train, validation, and test splits from VisDrone2019-DET [49]. The dataset features an average expression length of 9.0 words, with each expression referring to an average of 3.8 objects. Figure 1 gives several examples that relate to the three key features and challenges in RefDrone:
1) Multi-scale and small-scale target detection. Figure 3 presents an analysis of object scale distribution in RefDrone, with (a) showing the overall distribution and (b) offering a comparison with the gRefCOCO [22] dataset. In the RefDrone dataset, small-scale objects constitute 31% of all instances, while large-scale objects account for 14%. The detailed scale distribution in RefDrone demonstrates greater variance compared to gRefCOCO, highlighting the multi-scale challenges in RefDrone.
2) Multi-target and no-target samples. The RefDrone dataset includes 11,362 multi-target and 847 no-target expressions, with the number of referred targets ranging from 0 to 242. Figure 4 illustrates the target number distribution, revealing higher complexity in multi-target scenarios compared to gRefCOCO, where expressions typically refer to no more than two objects.
3) Complex environment with rich contextual expressions. The images present inherent complexity including diverse viewpoints, varying lighting conditions, and complex background environments. The referring expressions go beyond simple object attributes (e.g., color, category) and spatial relationships (e.g., ‘left’, ‘near’), incorporating rich object-object interactions (e.g., ‘the white trucks carrying livestock’) and object-environment interactions (e.g., ‘the white cars line up at the intersection’). Figure 5 presents word cloud visualizations of (a) complete expressions and (b) background terms, highlighting the diverse vocabulary used in descriptions.
3.4 Dataset comparison
Table 1 compares RefDrone with existing REC datasets. RefDrone stands out by its average number of referred objects and its use of LMM for expression generation. The expressions offer richer contextual details compared to template-based or human-annotated expressions. While RIS-CQ [13] employs LLM for expression generation, it lacks visual content during the generation process. This often results in expressions that, despite being linguistically complex, may be disconnected from visual content. Similarly, RSVG [47], although it focuses on small targets, falls short in terms of expression quality. In contrast, RefDrone comprehensively presents three challenges, establishing it as a challenging benchmark in REC tasks.
3.5 Evaluation metrics
We introduce instance-level metrics extending traditional REC metrics to address multi-target challenges. Previous benchmarks primarily focus on image-level metrics, which are sufficient for single-target or few-target samples. However, these metrics fail to handle expressions referring to multiple target objects—potentially scaling up to 100 instances per expression in our task. To address this limitation, we introduce instance-level metrics that provide a more granular evaluation of a model’s capability to identify individual target objects.
Instance-level metrics: Accinst. and F1inst. measure the accuracy of individual bounding box predictions. We compute the intersection over union (IoU) between the true bounding boxes and the predicted bounding boxes. An IoU $\geq$ 0.5 is considered a true positive (TP), otherwise it is a false positive (FP). Unmatched true bounding boxes are counted as false negatives (FN). For no-target samples, a prediction without any bounding box is a true negative (TN), otherwise it is a false positive (FP). We calculate:
$\text{Acc}_{inst.}=\frac{TP+TN}{TP+TN+FP+FN}$ (1)
and
[…]
Image-level metrics: Accimg. and F1img. evaluate the overall accuracy of predictions per image. An image is considered a true positive (TP) if all predictions match the true bounding boxes, otherwise it is a false positive (FP). For no-target samples, a prediction without any bounding box is a true negative (TN), otherwise it is a false positive (FP). We calculate Accimg. and F1img. using the same formulas as Accinst. and F1inst., respectively.
4 Proposed method
We introduce Number GroundingDINO (NGDINO) to address multi-target and no-target challenges based on GDINO [28]. Our insight is that utilizing number information of referred objects enhances model performance on these challenges. As illustrated in Figure 6, we introduce three components: (1) a number prediction head, (2) number-queries, and (3) a number cross-attention module.
Figure 6: Architecture of a single decoder layer in Number GroundingDINO. Key modifications from GDINO [28] (highlighted in yellow box) include: (1) number prediction head (FFN) to predict target number, (2) number-queries selected through the predicted number, (3) number cross-attention between selected number-queries and detection-queries.
4.1 NGDINO
Following GDINO, our model employs a dual-encoder-single-decoder architecture. The framework consists of an image backbone for visual feature extraction, a text backbone for text feature extraction, a feature enhancer for cross-modal fusion, a language-guided query selection module for detection-query initialization, and a cross-modality decoder for box refinement. Within decoder layers, detection-queries are fed into a self-attention layer, an image cross-attention layer to combine image features, a text cross-attention layer to combine text features, and an FFN layer. We mainly improve the decoder part, highlighted by the yellow box in Figure 6.
The number prediction head, structured similarly to the detection head with an FFN layer, predicts the number of referred objects from detection-queries. These detection-queries are used to predict the objects, containing number information. The number-queries are learnable embeddings designed to capture various numerical patterns. These queries are initialized randomly with dimensions (bs, length_nquery, ndim), where bs is the batch size, length_nquery represents the number of query lengths, and ndim denotes the feature dimension. The predicted number guides the selection of number-queries through the query selection in Algorithm 1. The selected number-queries have a fixed length of length_snquery.
We implement number cross-attention in parallel with self-attention to integrate the number information. In the number cross-attention, the selected number-queries serve as keys and values, while detection-queries serve as queries. The outputs are added to the self-attention features.
4.2 Loss function
[…]
""" Input: detection_queries:(bs,length_dquery,ndim) number_queries:(bs,length_nquery,ndim) Output: selected_number_queries:(bs,length_snquery,ndim) """ number_feat = FFN(detection_queries) # (bs, length_dquery, 1) number_feat = number_feat.mean(dim=1) # (bs, 1) number = sigmoid(number_feat) # (bs, 1) index = slice(length_snquery*number, length_snquery*(number+1)) selected_number_queries = number_queries[index] # (bs, length_snquery, ndim)
Algorithm 1 Query selection.
5 Experiments
5.1 Construction of RDAgent and NGDINO
RDAgent. A straightforward method is to adapt RDAgent for REC tasks. This adaptation automates the process by replacing the object input with predictions from an object detector and removing the human verification step. For a fair comparison, we employ Faster R-CNN [34] as the object detector, which is trained on the VisDrone2019-DET [49] dataset. The agent component utilizes GPT-4o [^1] https://openai.com/index/hello-gpt-4o/.
NGDINO. The NGDINO adopts a two-stage procedure to maintain training stability. First, we pre-train the number prediction head on the RefDrone dataset while initializing other components with GDINO [28] parameters. Then, we fine-tune the entire model.
5.2 Implementation details
We establish the benchmark with 13 representative models that can perform REC tasks, comprising 3 specialist models and 10 LMMs. The specialist models include MDETR [14], GLIP [19], and GDINO [28]. For LMMs, we evaluate Shikra [5], ONE-PEACE [38], SPHINX-v2 [21], MiniGPT-v2 [3], Ferret [44], Kosmos-2 [31], Griffon [48], Qwen-VL [1], CogVLM [41], and LLaVA [24]. Detailed model specifications are provided in the Supplementary Material.
Zero-shot evaluation details. For zero-shot evaluation, we use the original model checkpoints as provided in their respective papers. The implementations and checkpoints for GLIP and GDINO are obtained from the MMDetection [4].
Fine-tuning evaluation details. Our fine-tuning protocol maintains consistency across all experiments to ensure a fair comparison. We preserve the original learning strategies while excluding random crop augmentation due to its negative effect on position-sensitive samples. For LMMs, we employ the LoRA [11] fine-tuning strategy and follow the instruction tuning data structures. All fine-tuning experiments run for 5 epochs using 4 NVIDIA A100 GPUs.
5.3 Experimental results
Table 2: Experimental results of zero-shot baselines on RefDrone benchmark. Top groups: specialist models. Bottom groups: LMMs for REC tasks. GDINO-T and GDINO-B denote GroundingDINO [28] with Swin-Tiny and Swin-Base backbones. The best results in each group are denoted with bold.
Methods F1inst. Accinst. F1img. Accimg. MDETR [14] 8.42 4.41 2.99 1.63 GLIP [19] 5.46 3.84 9.20 8.54 GDINO-T [28] 1.18 1.84 3.94 6.35 GDINO-B [28] 1.97 2.23 6.43 7.58 Shikra [5] 0.80 0.52 2.26 1.60 ONE-PEACE [38] 1.02 0.51 2.64 1.34 SPHINX-v2 [21] 1.59 0.80 4.61 2.36 MiniGPT-v2 [3] 2.69 1.36 6.38 3.29 Ferret [44] 3.18 1.62 8.48 4.43 Kosmos-2 [31] 8.06 4.20 8.64 4.52 Griffon [48] 9.16 4.81 16.77 9.18 Qwen-VL [1] 10.91 5.77 18.32 10.08 CogVLM [41] 15.38 8.33 30.73 18.15
Zero-shot results. Table 2 presents the zero-shot evaluation results, assessing the models’ domain generalization capability. CogVLM [41] demonstrates state-of-the-art performance across multiple metrics. However, several advanced models, including Shikra [5], ONE-PEACE [38], and SPHINX-v2 [21], are limited to outputting only a single bounding box, due to constraints in their pre-training data or output strategy. This restriction significantly impacts their performance in multi-target scenarios.
Table 3: Experimental results of fine-tuning baselines on RefDrone benchmark. Top groups: specialist models. Bottom groups: LMMs for REC tasks.
Methods F1inst. Accinst. F1img. Accimg. MDETR [14] 32.60 19.49 19.17 10.81 GLIP [19] 24.23 14.86 16.92 13.29 GDINO-T [28] 30.50 19.02 29.65 21.07 NGDINO-T (Ours) 33.34 20.98 32.45 22.84 GDINO-B [28] 31.96 19.99 31.69 22.26 NGDINO-B (Ours) 34.44 21.76 34.01 23.89 MiniGPT-v2 [3] 4.97 2.74 13.56 8.97 LLaVA [24] 6.00 3.63 14.43 11.57 Qwen-VL [1] 14.14 7.61 20.10 11.17 RDAgent (Ours) 58.14 41.13 37.07 23.54
Fine-tuning results. Table 3 illustrates the fine-tuning performance across different methods. The specialist model MDETR [14] shows strong performance in instance-level metrics, achieving comparable results to GDINO-B [28], but struggles in image-level understanding. Conversely, LMMs like Qwen-VL [1] demonstrate superior image-level comprehension (F1img.: 20.10%, Accimg.: 11.17%) while struggling with instance-level tasks (F1inst.: 14.14%, Accinst.: 7.61%). This performance disparity can be attributed to LMMs’ strong global image understanding capabilities but limited effectiveness in detecting small objects due to input resolution constraints.
The proposed RDAgent achieves superior results compared to existing approaches, particularly in instance-level metrics, surpassing GDINO-B [28] by 26.18% in F1inst. and 21.14% in Accinst.. The relatively smaller gains in image-level metrics (5.38% in F1img. and 1.28% in Accimg.) highlight the importance of instance-level metrics. However, RDAgent’s multi-step pipeline and reliance on ChatGPT responses result in longer processing time compared to end-to-end approaches.
The proposed end-to-end method, NGDINO, consistently improves over the baseline GDINO [28] across both backbone architectures. Figure 7 shows the visualization results comparing NGDINO and baseline GDINO, demonstrating NGDINO’s effectiveness with multi-target samples.
Table 4: Experimental results on gRefCOCO [22] dataset. Asterisk (*) denotes results reported in the gRefCOCO paper.
Methods testA testB
Pr@0.5 $\uparrow$ N-acc. $\uparrow$ Pr@0.5 $\uparrow$ N-acc. $\uparrow$
MDETR* [14] 50.0 34.5 36.5 31.0
UNINEXT* [43] 46.4 49.3 42.9 48.2
GDINO-T [28] 45.69 79.02 44.83 76.69
NGDINO-T 46.05 83.17 45.57 78.08
To further validate the effectiveness of the proposed NGDINO, we conduct additional experiments on the gRefCOCO [22] dataset, which also includes multi-target and no-target samples (Table 4). NGDINO-T outperforms the baseline method GDINO-T, particularly in N-acc. metrics (by 4.15% in test A and 1.39% in test B), which evaluate accuracy on no-target samples. The limited improvements in Pr@0.5 metrics can be attributed to the relatively simple nature of multi-target samples in gRefCOCO, which typically contain only two objects per sample.
Figure 7: Visualization results comparing NGDINO and GDINO [28], where red boxes indicate true positives and yellow boxes denote false negatives.
Figure 8: Failure cases of NGDINO on RefDrone dataset. Red, green, and yellow boxes indicate true positives, false positives, and false negatives, respectively.
Table 5: Ablation study on RefDrone dataset for the effectiveness of each component in NGDINO, with Swin-Tiny backbone. NPH denotes number prediction head. NCA denotes number cross-attention.
NPH NCA Acc_$\_inst.$ Acc_$\_img.$ FPS 19.02 21.07 13.5 19.09 (0.07$\uparrow$) 21.18 (0.11$\uparrow$) 12.8 (0.7$\downarrow$) 19.51 (0.49$\uparrow$) 21.71 (0.53$\uparrow$) 12.9 (0.6$\downarrow$) 20.98 (1.96$\uparrow$) 22.84 (1.77$\uparrow$) 12.3 (1.2$\downarrow$)
Table 6: Ablation study for the impact of varying selected number query length. Params indicates additional parameters introduced.
Length F1inst. Accinst. F1img. Accimg. Params 1 31.63 19.80 31.23 22.12 1.58M 10 33.34 20.98 32.45 22.84 1.65M 100 32.06 20.10 32.23 22.81 2.34M
Table 7: Ablation study for the effectiveness of RDAgent.
Methods F1inst. Accinst. F1img. Accimg. ReCLIP [36] 24.62 14.04 11.58 6.15 GPT4-o 52.38 35.65 35.50 22.38 RDAgent 58.14 41.13 37.07 23.54
5.4 Ablation studies
Ablations of NGDINO components. Table 5 presents the analysis of each component in NGDINO. The number prediction head alone shows a slight increase in Accinst. from 19.02% to 19.09%, indicating minimal impact of this auxiliary head on the referring tasks. When introducing the number cross-attention without number selection, we observe a more substantial improvement, with Accinst. increasing from 19.02% to 19.51%. This improvement is attributed to the additional parameters introduced in the decoder. The significant improvement is achieved through the combination of both the number prediction task and number cross-attention components. This contribution increases Accinst. from 19.02% to 20.98%. While the additional computational cost results in a minor decrease in inference speed (FPS), this trade-off yields a 1.96% increase in Accinst..
Ablations of query length. Table 6 analyzes the impact of varying the selected number query length. Utilizing a minimal query length of 1 lacks the capacity to capture complex numerical information. Conversely, extending the query length to 100 increases parameter count and introduces optimization difficulties, thereby adversely affecting performance. Through these experiments, we determine that a query length of 10 provides an optimal trade-off.
Number prediction head performance. The number prediction head achieves an overall number prediction accuracy of 53.5%. During the bounding box prediction stage, this accuracy is observed to be 22.84%. This performance gap demonstrates the effectiveness of number prediction head. Besides, the number prediction achieves a mean absolute error (MAE) of 0.51, suggesting high precision as predictions closely align with ground truth values.
Ablations of RDAgent. Table 7 analyzes the effectiveness of the proposed RDAgent. The GPT4-o baseline is established using only Step 3 of RDAgent. We also compare against ReCLIP [36], a two-stage instance ranking method using CLIP [33] for REC tasks. All experiments utilize Faster-RCNN [34] for object detection. Results show that RDAgent consistently outperforms GPT4-o and ReCLIP across all metrics. However, the performance is partially limited by the Faster-RCNN object detector, which achieves only 18.0 mAP on the VisDrone2019-DET [49] dataset.
5.5 Limitations
While NGDINO addresses the multi-target and no-target challenges, several challenges remain. Figure 8 shows typical NGDINO failure cases, highlighting challenges in the RefDrone dataset. The examples show rich contextual expressions, challenging backgrounds, and small-scale object detection. These challenging cases reflect real-world applications and highlight the areas for future improvement.
6 Conclusion
In this work, we introduce RefDrone, a challenging benchmark specifically designed for referring expression comprehension in drone scenes. The dataset is constructed using the proposed RDAgent, a semi-auto annotation framework leveraging a multi-agent system. Furthermore, we develop NGDINO to address the multi-target and no-target challenges in the RefDrone dataset. In the future, we aim to further enhance NGDINO to address additional challenges presented by RefDrone. We also plan to expand the benchmark to include referring expression segmentation and referring expression tracking tasks.
Supplementary Material
We provide supplementary material for further analysis, organized as follows:
  1. 1.
    Color attribute analysis of RefDrone. (Appendix A)
  2. 2.
    Details of baseline method. (Appendix B)
  3. 3.
    Additional dataset examples. (Appendix C)
  4. 4.
    Resutlts w.r.t different object scales. (Appendix D)
  5. 5.
    Prompts and examples in RDAgent. (Appendix E)
Appendix A Dataset color attribute analysis
In the RefDrone dataset, color attributes are present in 69% expressions. We initially use the HSV color space to define six primary colors: white, black, red, blue, green, and yellow. During the annotation process, RDAgent expands the color set by adding six more colors: orange, pink, grey, purple, brown, and silver. The distribution of these color terms across the expressions is illustrated in Figure 9.
Figure 9: Distribution of color terms in RefDrone expressions.
Appendix B Details of baseline methods
The details for each baseline method:
  • Specialist models:
    • MDETR: ResNet-101 with BERT-Base,
      pretrained on Flickr30k, RefCOCO/+/g, VG.
    • GLIP: Swin-Tiny with BERT-Base,
      pretrained on Object365.
    • GDINO-T: Swin-Tiny with BERT-Base,
      pretrained on Object365, GoldG, GRIT, V3Det.
    • GDINO-B: Swin-Base with BERT-Base,
      pretrained on Object365, GoldG, V3Det.
  • Large multimodal models:
    • Kosmos-2: 1.6B parameters.
    • ONE-PEACE: 4B parameters, visual grounding API.
    • Shikra: 7B parameters, delta-v1 version.
    • MiniGPT-v2: 7B parameters.
    • LLaVA: 7B parameters, v1.5 version.
    • Qwen-VL: 7B parameters.
    • Ferret: 7B parameters.
    • CogVLM: 7B parameters, grounding-specific version.
    • SPHINX-v2, Griffon: 13B parameters.
Appendix C Dataset examples
To provide a comprehensive understanding of our RefDrone dataset, we present representative examples in Figure 10. These samples demonstrate the three key challenges in our dataset, highlighting its real-world applicability.
Appendix D The results of different object scales.
We add small/medium/large instance accuracy metrics (ACCs, ACCm, ACCl) of RefDrone dataset in Table 8.
Table 8: Fine-tuned results w.r.t. different object scales.
Methods Accs Accm Accl Methods Accs Accm Accl GLIP 2.65 18.28 27.29 LLaVA 4.18 7.26 11.99 GDINO-T 8.56 20.66 33.99 Qwen-VL 0.73 10.60 30.18 NGDINO-T 10.84 23.38 40.13 RDAgent 32.54 47.97 47.17
Appendix E Prompts and examples for RDAgent
In this section, we provide the prompts and examples employed in RDAgent. Table 9 presents the prompt construction process for expression generation (Step 3), which includes the system prompt and few-shot in-context learning examples. One in-context learning example is illustrated in Table 10. The system prompts used for each step are detailed in Table 11. Additionally, the system prompts for the feedback mechanism are presented in Table 12.
Figure 10: Dataset samples from RefDrone.
Table 9: Illustration of RDAgent’s prompt construction for expression generation (Step 3). Few-shot in-context-learning examples are from fewshot_samples. A representative example is shown in Figure 10
messages = [ {"role":"system", "content": f"""As an AI visual assistant, your role involves analyzing a single image. You are supplied with three sentences that caption the image, along with additional data about specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. When coordinate x tends to 0, the object nears the left side of the image, shifting towards the right as coordinate x approaches 1. When coordinate y tends to 0, the object nears the top of the image, shifting towards the bottom as coordinate y approaches 1. Your task is to classify the provided objects based on various characteristics, while also substantiating your classification. This classification should be thoroughly justified, with criteria including but not limited to relationships or relative locations of objects. To refer to a specific object, use the provided coordinates directly. Base your classification justifications on direct observations from the image, avoiding any hypothesizing or assumptions."""} ] for  sample in fewshot_samples:
       messages.append({"role":"user", "content":sample[‘context’]})
messages.append({"role":"assistant", "content":sample[‘response’]} )
[…]
Table 10: One example to illustrate the input and output in RDAgent Step 3 (expression generation). The top block shows the input contexts such as image, object, caption and color attributed. The bottom block shows the generated annotations.
Context type 1: Captions
The image shows an aerial view of a busy urban street with multiple lanes of traffic, flanked by parking areas and buildings. A median strip with greenery runs down the center, dividing the opposite flows of vehicles. On either side of the road, there are commercial and residential structures. Parking lanes on both sides of the main road are filled with parked vehicles.
[…]
Response : The blue cars park on the right parking lot. Coordinates: [0.628, 0.478], [0.646, 0.692], [0.649, 0.727] The white cars park on both sides of the main road in a line. Coordinates: [0.431, 0.051], [0.431, 0.15], [0.431, 0.101], [0.434, 0.256], [0.439, 0.523], [0.443, 0.813], [0.445, 0.902], [0.446, 0.982], [0.235, 0.43], [0.232, 0.486], [0.223, 0.674], [0.215, 0.904]
Table 11: System prompts employed at different stages of RDAgent, replacing the default system messages shown in Table 9
System prompt: Step 1 scene understanding
You are an AI visual assistant that specializes in providing clear and accurate descriptions of images without any ambiguity or uncertainty. Your descriptions should focus solely on the content of the image itself and avoid mentioning any location-specific details such as regions or countries where the image might have been captured.
System prompt: Step 2 color categorization
As an AI visual assistant, your role involves analyzing a single image.
You are supplied with the specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. Your task is to assess whether the given colors of specific objects match their appearance in the image. Respond with ”Yes” when the colors are appropriate. In cases where the colors are deemed inappropriate, respond with a concise ”No.” System prompt: Step 3 expression generation
As an AI visual assistant, your role involves analyzing a single image. You are supplied with three sentences that caption the image, along with additional data about specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. When coordinate x tends to 0, the object nears the left side of the image, shifting towards the right as coordinate x approaches 1. When coordinate y tends to 0, the object nears the top of the image, shifting towards the bottom as coordinate y approaches 1.
Your task is to classify the provided objects based on various characteristics, while also substantiating your classification. This classification should be thoroughly justified, with criteria including but not limited to relationships or relative locations of objects. To refer to a specific object, use the provided coordinates directly. Base your classification justifications on direct observations from the image, avoiding any hypothesizing or assumptions. System prompt: Step 4 quality evaluation
As an AI visual assistant, your role involves analyzing a single image. You are supplied with three sentences that describe the image, along with additional data about specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. When coordinate x tends to 0, the object nears the left side of the image, shifting towards the right as coordinate x approaches 1. When coordinate y tends to 0, the object nears the top of the image, shifting towards the bottom as coordinate y approaches 1. Besides, you are supplied with the description of the objects and their corresponding attributes.
Your task is to confirm whether the description exclusively relates to the described objects without including any others in the visual. Respond ”yes” if it matches, or ”no” with an explanation if it does not.
Table 12: RDAgent system prompts for feedback mechanism. Differences from Table 11 are highlighted
System prompt: Step 2 color categorization with feedback mechanism
As an AI visual assistant, your role involves analyzing a single image.
You are supplied with the specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. Your task is to assess whether the given colors of specific objects match their appearance in the image. Respond with ”Yes” when the colors are appropriate. In cases where the colors are deemed inappropriate, respond with a concise ”No.” System prompt: Step 3 expression generation with feedback mechanism
As an AI visual assistant, your role involves analyzing a single image. You are supplied with three sentences that caption the image, along with additional data about specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. When coordinate x tends to 0, the object nears the left side of the image, shifting towards the right as coordinate x approaches 1. When coordinate y tends to 0, the object nears the top of the image, shifting towards the bottom as coordinate y approaches 1. You are also provided with descriptions and the objects that initially failed to match, along with the reasons for the discrepancies.
Your task is to revise both the description and the corresponding objects to correct these mismatches based on the provided reasons. Ensure that the revised description accurately matches the corresponding objects depicted in the visual content.
System prompt: Step 4 quality evaluation with feedback mechanism
As an AI visual assistant, your role involves analyzing a single image. You are supplied with three sentences that describe the image, along with additional data about specific attributes of objects within the image. This can include information about categories, colors, and precise coordinates. Such coordinates, represented as floating-point numbers that range from 0 to 1, are shared as center points, denoted as (x, y), identifying the center x and y. When coordinate x tends to 0, the object nears the left side of the image, shifting towards the right as coordinate x approaches 1. When coordinate y tends to 0, the object nears the top of the image, shifting towards the bottom as coordinate y approaches 1. Besides, you are supplied with the description and the objects that initially failed to match.
Your task is to provide detailed reasoning for unsuccessful object matches.